OpenStudy (anonymous):

subtract one and one third from two and three fourth first write the fractions with the 12 as the common denominator

OpenStudy (anonymous):

12/12, 5/12

OpenStudy (anonymous):

one and one third = 4/3 = 16/12 two and three fourths = 11/4 = 33/12 33 16 17 -- - --- = ----- 12 12 12